# Revisiting Stochastic Gravitational-wave Background in the Strong Signal Case

### v1

@@ Abstract
Weak-signal limit is often used in estimating stochastic gravitational-wave background (SGWB) intensities. This approximation fails and the signal-to-noise ratio (SNR) can be much weaker when background signals are loud compared to the detector noise. In this work, we highlight this limitation for the SGWB detection using space-borne detector networks. For the TianQin + LISA network, the SNR estimated under the weak-signal limit might be off by as large as an order of magnitude. Contour plots of SNR over the parameter spaces are also presented to indicate regions susceptible to this discrepancy. Our results suggest that DA and DB type extragalactic double white dwarfs may yield an SGWB with SNR surpassing 100 after 1 year of operation in the weak-signal-limit scenario, with a redshift-independent merger rate of about $500\,\,{\rm Mpc^{-3}\,Myr^{-1}}$. In fact, this value falls significantly below the necessary threshold. Similar influences arise for first-order phase transitions, yet pinning down parameter regions remains formidable due to model uncertainties.
@@ I Introduction
Space-borne GW detectors, like TianQin Luo et al. (2016), Laser interferometry Space Antenna (LISAAmaro-Seoane et al. (2017), are typically composed of three satellites arranged in an equilateral triangle formation. However, the unequal armlengths resulting from detector motion present a challenge in canceling laser noise, leading to the requirement of time delay interferometry (TDI) technique Tinto and Armstrong (1999); Tinto and Dhurandhar (2021). Among the various TDI combinations, the unequal-arm Michelson with the noise-orthogonal basis (A,E,T) is commonly employed. The A/E channels are specifically tailored for detecting GW, providing optimized detection sensitivity Vallisneri et al. (2008). The T channel exhibits a substantially lower sensitivity to GW compared to the A/E channels, making it more suitable for monitoring detector noise rather than efficiently detecting GW Hogan and Bender (2001).
@@
One often relies on the auto- or cross-correlation of the channel data to derive the power spectral density (PSD) of SGWB. The null-channel method utilizes the auto-correlation of the null channel to effectively track and isolate noises, allowing for the identification of the SGWB component within a single space-borne detector Tinto et al. (2001); Hogan and Bender (2001); Adams and Cornish (2014); Boileau et al. (2021a); Muratore et al. (2022); Cheng et al. (2022). While in scenarios where a network of space-borne detectors is deployed, each responding to a common SGWB but with uncorrelated noises, distinguishing between the SGWB and detector noise can be achieved through cross-correlating the channel data from two or more detectors Hellings and Downs (1983); Michelson (1987); Christensen (1992); Flanagan (1993); Allen and Romano (1999); Ungarelli and Vecchio (2001); Abbott et al. (2018); Hu et al. (2024).
@@
Following the null-channel and cross-correlation methods, the expectation of correlation measurement is primarily determined by the PSD of SGWB and the overlap reduction function (ORF), which quantifies the reduction in the correlation to an SGWB resulting from the relative separation and orientation of the channels. On the other hand, the variance of correlation measurement is solely influenced by the noise PSD, given the SGWB is significantly lower than the detector noise (weak-signal limit). Based on the expectation and variance, the detection signal-to-noise ratio (SNR) can be calculated, taking into account factors such as correlation time accumulation and frequency band integration. Building upon the SNR for the weak-signal limit, the power-law integrated sensitivity (PLIS) and peak-integrated sensitivity (PIS) sensitivity curve is derived and commonly used to showcase the detection capability for the SGWB Thrane and Romano (2013); Schmitz (2021). Additionally, inversely solving the SNR can be employed to constrain the detectable regions of source parameters, as exemplified by SNR contour plots Caprini et al. (2016); Kuroyanagi et al. (2018); Caprini et al. (2020).
@@
While the weak-signal limit is applicable in most scenarios, there may exist SGWB with intensities matching or exceeding the detector noise Amaro-Seoane et al. (2012); Pan and Yang (2020); Huang et al. (2020); Sharma and Harms (2020); Zhou et al. (2023); Wu and Li (2023); Torres-Orjuela et al. (2023); Mentasti et al. (2023); Song et al. (2024). Therefore, Allen et al. extended the SNR to handle arbitrarily large SGWB, specifically tailored for an ORF-invariant channel pair Allen and Romano (1999); Cornish (2002); Kudoh et al. (2006). Their findings suggested that the existence of SGWBs can impact the variance of correlation measurements, potentially resulting in an overestimation of detection capability. This SNR estimator has been applied in a series of subsequent studies Wang and Han (2021); Brzeminski et al. (2022); Liang et al. (2023a); Hu et al. (2024). However, to the best of our knowledge, it has not been utilized for constructing PLIS and PIS sensitivity curves, nor in generating SNR contour plots.
@@
In this paper, we further develop a more robust SNR estimation, adaptable to multiple channel pairs and time-varying ORF. By incorporating the correction of the SNR estimation, we reconstruct the PLIS and PIS sensitivity curves. Focusing on TianQin, the TianQin I+II network, and the TianQin + LISA network, we identify gaps in sensitivity curves between the scenarios of approximation and non-approximation. Additionally, to illustrate the gaps in constraining detectable parameter regions between these two scenarios, we employ the SNR contour plot.
@@ II Space mission
As shown in Fig. 1, TianQin is a space-borne GW detector comprising three identical satellites in Earth’s orbit. These satellites will have a 3.64-day orbital period and will be positioned at a radius of approximately $10^{5}\,\,{\rm km}$. When operational, they will form an equilateral triangle with an arm length ($L_{\rm TQ}$) of about $1.7\times 10^{5}\,\,{\rm km}$. TianQin will operate on a “three months on + three months off” schedule, with a proposed second mission, TianQin II, aimed at filling detection gaps Ye et al. (2019). The orbital planes of TianQin and TianQin II are arranged perpendicular to each other, creating the TianQin I+II network. The nominal work scheme leads to no overlap in the operation times of TianQin and TianQin II; we instead consider a scenario where the observation time can be extended and correspond to an enhanced operation mode of “four months on + two months off”.
@@ II.2 TDI channel
Due to the detector motion, maintaining an equilateral shape for a space-borne detector is challenging, making it difficult to implement an equal-arm Michelson to cancel laser noise. In response to this challenge, the TDI combination has been proposed Tinto and Armstrong (1999). The first-generation TDI is initially designed for static detectors and includes components such as the Sagnac, unequal-arm Michelson, Relay, Monitor, and Beacon Tinto and Dhurandhar (2021). This paper focuses on the unequal-arm Michelson (X,Y,Z), which can be established using each satellite and its adjacent link. Furthermore, the corresponding noise-orthogonal unequal-arm Michelson can be constructed by Vallisneri et al. (2008)
@@ III.1 Stochastic background
To begin, we constrain our focus on the Gaussian-stationary, unpolarized, and isotropic SGWB. Therefore, it is reasonable to assume that the SGWB has a zero-mean Fourier amplitude $\widetilde{h}_{P}(f,\hat{k})$, with the time-independent variance:
@@
Here the factor $1/4$ stems from both the definition of one-side PSD and the average of polarization. The PSD of SGWB incorporates contributions from all directions in the sky, as reflected in the all-sky integral of the $\mathscr{P}_{\rm h}$:
@@
To further characterize the distribution of energy across different frequencies in the SGWB, a commonly used quantity is the dimensionless energy spectrum density $\Omega_{\rm gw}$. This quantity serves to capture the ratio of GW energy density ${\rm d}\rho_{\rm gw}$ within a specific frequency range [$f$,$f+{\rm d}f$] to the critical energy density $\rho_{\rm gw}$, exhibiting a direct relationship with $S_{\rm h}$ Allen (1996); Thrane and Romano (2013):
@@ III.3 Detection sensitivity
The channel data $s_{I}$ is typically a combination of the SGWB signal $h_{I}$ and the channel noise $n_{I}$. As the SGWB signal and the channel noise are expected to be uncorrelated, in order to extract the SGWB signal from the data, a correlation measurement needs to be constructed as follows:
$S_{IJ}=\int_{t_{0}-T/2}^{t_{0}+T/2}{\rm d}t\int_{t_{0}-T/2}^{t_{0}+T/2}{\rm d}t^{\prime}\left[s_{I}(t)s_{J}(t^{\prime})-\langle n_{I}(t)n_{J}(t^{\prime})\rangle\right],$ (12)
where the noise PSD is determined by
@@
By monitoring the noise using the null-channel method Tinto et al. (2001); Hogan and Bender (2001), or achieving a zero-value noise PSD through the cross-correlation method Christensen (1992); Flanagan (1993), it is possible to identify the SGWB signal. These methods involve different channels from the same and different detectors. To streamline subsequent analysis, we designate the detectors as $\mathcal{A}$ and $\mathcal{B}$. Assuming that the employed channels in each detector share the same auto-correlation ORF $\mathcal{R}_{\mathcal{A}}$ and noise PSD, the total SNR in the weak-signal limit Seto (2020); Liang et al. (2022)
@@
the SNR threshold $\rho_{\rm thr}$ can be translated into the threshold of SGWB strength $\Omega_{0}$ based on Eq. (14):
@@ IV Beyond weak-signal limit
Assuming that all other conditions specified in Eq. (12) hold true, except for the weak-signal limit. Then for all channel pairs $\{IJ\}$ (or equivalently, $\{MN\}$), the expectation and variance of the correlation measurement
@@
where the filter function $\widetilde{Q}$ is arbitrary. Furthermore, for the last line of Eq. (25) to be valid, it is necessary that the ORF for cross-correlation is significantly lower than that for auto-correlation, leading to
@@
$\rho_{\rm opt}^{2}=\frac{2}{1+\delta_{\mathcal{A}\mathcal{B}}}\frac{\big{(}\widetilde{Q}_{IJ}(f,t_{0}),\frac{f^{3}\Gamma^{*}_{IJ}(f,t_{0})\Omega_{\rm gw}(f)}{D_{IJ}(f)}\big{)}^{2}}{\big{(}\widetilde{Q}_{IJ}(f,t_{0}),\widetilde{Q}_{IJ}(f,t_{0})\big{)}}.$ (29)
@@
$\widetilde{Q}_{IJ}(f,t_{0})=\lambda\frac{f^{3}\Gamma^{*}_{IJ}(f,t_{0})\Omega_{\rm gw}(f)}{D_{IJ}(f)},$ (30)
@@
In contrast to Eq. (14), Eq. (31) suggests that the SGWB detection can suffer negative impacts from the SGWB itself, especially in cases of loud SGWB. To visualize this impact, plotting sensitivity curves is beneficial.
Although a direct derivation of $\Omega_{0}$ and $\Omega_{\rm PIS}$ directly from Eq. (18) and Eq. (22) based on Eq. (31) to generate the sensitivity curves proves challenging, these values can still be determined by inverting the SNR formula [^1] In this work, we utilize the Newton-Raphson method.. Fig. 3 and Fig. 4 illustrate the PLIS and PIS curves correspondingly, with solid lines in red, green, and blue denoting TianQin, the TianQin I+II network, and the TianQin + LISA network  [^2] Note that for the TianQin detector, both auto-correlations within the same channel and cross-correlations between different channels are considered. However, in the case of detector networks, only the cross-correlation between channels from different detectors is taken into consideration., surpassing the weak-signal limit. Dashed lines including the weak-signal limit are provided for comparison. Note that the spectral function for the PIS curves is determined by
@@ V Case study
Furthermore, it has been highlighted that the SGWB originating from extragalactic double white dwarfs may exhibit significant intensity levels, leading to high SNR Liang et al. (2022). Without loss of generality, the energy spectrum density of SGWB can be described as the cumulative contribution from each binary system Abbott et al. (2016, 2018):
@@
where the population parameter $\bm{\theta}$ primarily encompasses the component masses of the system. In this paper, we focus on the DA and DB white dwarfs, for which the mass distribution can be modeled by a Gaussian distribution with mean $\mu=0.663\,\,M_{\odot}$ and standard deviation $\sigma=0.177\,\,M_{\odot}$ Rosado (2011). The merger rate is directly linked to the redshift and can be expressed as follows Abbott et al. (2023):
@@
with the local merger rate $R_{0}$ and index $\kappa$. As suggested in Ref. Rosado (2011), the $R_{0}$ is assumed to be within the range of 20 to $500\,\,{\rm Mpc^{-3}\,yr^{-1}}$. The energy spectrum emitted in the source frame, denoted by ${\rm d}E_{\rm gw}/{\rm d}f_{\rm s}$, is characterized by the frequency in the source frame $f_{\rm s}=f(1+z)$ Zhu et al. (2013). Given that the radius of a white dwarf significantly exceeds its Schwarzschild radius, it is reasonable to assume that below the cut-off frequency where both stars come into contact, the binary systems consistently remain in the inspiral phase.
@@
Regarding the first-order PTs, we adopt the model mentioned in our previous study Liang et al. (2022), where the energy spectrum density is formulated as
@@
$\tilde{\Omega}_{i}(\bm{\theta})=1.67\times 10^{-5}\left(\frac{100}{g_{*}(T_{*})}\right)^{1/3}\left(\frac{H_{*}}{\beta}\right)^{p}\left(K_{i}(\alpha)\frac{\alpha}{1+\alpha}\right)^{q}\tilde{\Delta}_{i*}(v_{w}),$ (38)
@@
Figure. 5 illustrates the SGWBs from various sources: the purple line represents the Galactic foreground with a one-year operation time, the gray region corresponds to the extragalactic background associated with the local merger rate ranging from 20 to $500\,\,{\rm Mpc^{-3}\,yr^{-1}}$, and the orange line signifies the cosmological background determined by the parameters $\{\alpha,\alpha_{\infty},v_{w},\beta/H_{*},T_{*}\}=\{2,1,0.95,50,10^{3}\,{\rm{GeV}}\}$. To facilitate comparison between the SGWB intensity and noise levels, the noise energy spectrum density of TianQin, the TianQin I+II network, and the TianQin + LISA network are shown in red, green, and blue lines, respectively. While the extragalactic background can near the detector noise level, both the Galactic foreground and cosmological background can exceed the noise. Taking the Galactic foreground as an example, Table. 3 presents the detection SNR within different detector configurations. It also provides the SNR under weak-signal conditions for the sake of comparison. In the weak-signal limit, the TianQin + LISA network achieves the highest SNR for detecting the Galactic foreground. As moving beyond the weak-signal limit, the SNR drops to approximately 20% for the TianQin + LISA network, whereas it remains around 60% for TianQin and the TianQin I+II network. Thus, TianQin demonstrates superior performance in terms of detecting Galactic foreground.
@@
In Fig. 6, we present SNR contour plots for the extragalactic background, showcasing scenarios with and without the weak-signal limit in the left and right panels for comparison. Not surprisingly, an SNR threshold of 100 leads to significant variations in the detectable parameter regions between the two scenarios, especially for the TianQin + LISA network. Failure to achieve an SNR above 100 within one year of operation would shift the lower limit for $\kappa$ from around 0 to 2.5, indicating a transition from a scenario with the weak-signal limit to one without it. This shift suggests that the adverse effects of loud SGWB self-generation on detection reduce the region of detectable parameters to a lesser extent than initially anticipated. On the other hand, assuming a constant merger rate with redshift ($\kappa=0$), achieving an SNR exceeding 100 after 1 year of operation, under the weak-signal limit, would estimate the merger rate for EDWDs to be around $500\,\,{\rm Mpc^{-3}\,yr^{-1}}$. However, this value falls considerably short of meeting the actual detection requirements.
A similar diminishing effect of parameter constrains can be observed in Fig. 7, where we set the parameters $\{\alpha,\alpha_{\infty},v_{w}\}=\{2,1,0.95\}$ for the first-order PTs. If the SNR surpasses 100 after 1 year of operation, the temperature $T_{*}$ is expected to be significantly greater than $10\,\,{\rm{GeV}}$, while the PT duration $\beta/H_{*}$ should be less than $10^{2}$. By leveraging the weak-signal limit, the detectable parameter region for these two variables can significantly expand, potentially leading to misestimations. It is important to highlight that, due to the wide range of variability in fixed parameters, effectively constraining the parameters of PTs remains a challenging task.
Figure 6: SNR contour plots for the SGWB resulting from EDWD. The left-hand and right hand panels refer to the scenarios with and without weak-signal limit, considering a one-year operation time.

### v2

# Revisiting Stochastic Gravitational Wave Background in the Strong Signal Case

@@ Abstract
Weak-signal limit is often used in estimating stochastic gravitational-wave background (SGWB) intensities. This approximation fails and the signal-to-noise ratio (SNR) can be much weaker when background signals are loud compared to the detector noise. In this work, we find that this issue is especially significant when dealing with networks of detectors that are widely separated. For the TianQin + LISA network, the SNR estimated under the weak-signal limit might be off by as large as an order of magnitude. Contour plots of SNR over the parameter spaces are also presented to indicate regions that will be affected by this correction. Our results suggest that DA and DB type extragalactic double white dwarfs may yield an SGWB with SNR surpassing 100 after 1 year of operation in weak-signal limit scenario, with a redshift-independent merger rate of about $500\,\,{\rm Mpc^{-3}\,Myr^{-1}}$. In fact, this value falls significantly below the necessary threshold. Similar influences arise for first-order phase transitions, yet pinning down the detectable parameters remains formidable due to model uncertainties.
@@ I Introduction
Space-borne GW detectors, like TianQin Luo et al. (2016), Laser Interferometer Space Antenna (LISAAmaro-Seoane et al. (2017), are typically composed of three satellites arranged in an equilateral triangle formation. However, the unequal armlengths resulting from detector motion present a challenge in canceling laser noise, leading to the requirement of time delay interferometry (TDI) technique Tinto and Armstrong (1999); Tinto and Dhurandhar (2021). Among the various TDI combinations, the unequal-arm Michelson with the noise-orthogonal basis (A,E,T) is commonly employed. The A/E channels are specifically tailored for detecting GW, providing optimized detection sensitivity Vallisneri et al. (2008). The T channel exhibits a substantially lower sensitivity to GW compared to the A/E channels, making it more suitable for monitoring detector noise rather than efficiently detecting GW Hogan and Bender (2001).
@@
One often relies on the auto- or cross-correlation of the channel data to derive the power spectral density (PSD) of SGWB. The null-channel method utilizes the auto-correlation of the null channel to effectively track and isolate noises, allowing for the identification of the SGWB component within a single space-borne detector Tinto et al. (2001); Hogan and Bender (2001); Adams and Cornish (2014); Smith and Caldwell (2019); Boileau et al. (2021a); Muratore et al. (2022); Cheng et al. (2022). While in scenarios where a network of space-borne detectors is deployed, each responding to a common SGWB but with uncorrelated noises, distinguishing between the SGWB and detector noise can be achieved through cross-correlating the channel data from two or more detectors Hellings and Downs (1983); Michelson (1987); Christensen (1992); Flanagan (1993); Allen and Romano (1999); Ungarelli and Vecchio (2001); Abbott et al. (2018); Hu et al. (2024).
@@
Following the null-channel and cross-correlation methods, the expectation of detection measurement is primarily determined by the PSD of SGWB and the overlap reduction function (ORF), which quantifies the reduction in the correlation to an SGWB resulting from the relative separation and orientation of the channels. On the other hand, the variance of detection measurement is solely influenced by the noise PSD, given the SGWB is significantly lower than the detector noise (weak-signal limit). Based on the expectation and variance, the detection signal-to-noise ratio (SNR) can be calculated, taking into account factors such as correlation time accumulation and frequency band integration. Building upon the SNR for the weak-signal limit, the power-law integrated sensitivity (PLIS) and peak-integrated sensitivity (PIS) sensitivity curve is derived and commonly used to showcase the detection capability for the SGWB Thrane and Romano (2013); Schmitz (2021). Additionally, inversely solving the SNR can be employed to constrain the detectable regions of source parameters, as exemplified by SNR contour plots Caprini et al. (2016); Kuroyanagi et al. (2018); Caprini et al. (2020).
@@
While the weak-signal limit is applicable in most scenarios, there may exist SGWB with intensities matching or exceeding the detector noise  Amaro-Seoane et al. (2012); Pan and Yang (2020); Huang et al. (2020); Sharma and Harms (2020); Zhou et al. (2023); Wu and Li (2023); Torres-Orjuela et al. (2023); Mentasti et al. (2023); Song et al. (2024). Therefore, Allen et al. extended the SNR to handle arbitrarily large SGWB, specifically tailored for an ORF-invariant channel pair  Allen and Romano (1999); Cornish (2002); Kudoh et al. (2006). They highlighted that the existence of SGWBs can impact the variance of detection measurements, potentially resulting in an overestimation of detection capability Allen and Romano (1999); Cornish (2002); Kudoh et al. (2006). This SNR estimator has been applied in a series of subsequent studies Wang and Han (2021); Brzeminski et al. (2022); Liang et al. (2023a); Hu et al. (2024). However, to the best of our knowledge, it has not been utilized for constructing PLIS and PIS sensitivity curves, nor in generating SNR contour plots.
@@
In this paper, we further develop a more robust SNR estimation, adaptable to multiple channel pairs and time-varying ORF. By incorporating the correction of the SNR estimation, we revise the PLIS and PIS sensitivity curves. Focusing on TianQin, the TianQin I+II network, and the TianQin + LISA network, we identify gaps in sensitivity curves between the scenarios of approximation and non-approximation. Additionally, to illustrate the gaps in constraining detectable parameter regions between these two scenarios, we employ the SNR contour plot.
@@ II Space mission
As shown in Fig. 1, TianQin is a space-borne GW detector comprising three identical satellites in Earth’s orbit. These satellites will have a 3.64-day orbital period and will be positioned at a radius of approximately $10^{5}\,\,{\rm km}$. When operational, they will form an equilateral triangle with an arm length ($L_{\rm TQ}$) of about $1.7\times 10^{5}\,\,{\rm km}$. TianQin will operate on a “three months on + three months off” schedule, with a proposed second mission, TianQin II, aimed at filling detection gaps Ye et al. (2019). The orbital planes of TianQin and TianQin II are arranged perpendicular to each other, creating the TianQin I+II network. The nominal work scheme leads to no overlap in the operation times of TianQin and TianQin II; we instead consider a scenario where the observation time can be extended and correspond to an enhanced operational mode of “four months on + two months off”.
@@ II.2 TDI channel
Due to the detector motion, maintaining an equilateral shape for a space-borne detector is challenging, making it difficult to implement an equal-arm Michelson to cancel laser noise. In response to this challenge, the TDI combination has been proposed Tinto and Armstrong (1999). The first-generation TDI is initially designed for static detectors and includes components such as the Sagnac, unequal-arm Michelson, Relay, Monitor, and Beacon Tinto and Dhurandhar (2021). This paper focuses on the unequal-arm Michelson (X,Y,Z), which can be established using each satellite and its adjacent link. Furthermore, the corresponding noise-orthogonal bases can be constructed by Vallisneri et al. (2008)
@@ III.1 Stochastic background
To begin, we narrow our focus to the Gaussian-stationary, unpolarized, and isotropic SGWB. Therefore, it is reasonable to assume that the SGWB possesses a zero-mean Fourier amplitude $\widetilde{h}_{P}(f,\hat{k})$, with the direction-dependent PSD $\mathscr{P}_{\rm h}$:
@@
Here the factor $1/4$ stems from both the definition of one-side PSD and the average of polarization, $\delta_{ij}$ denotes the Kronecker delta. The PSD of SGWB incorporates contributions from all directions in the sky, as reflected in the all-sky integral of the $\mathscr{P}_{\rm h}$:
@@
To further characterize the distribution of energy across different frequencies in the SGWB, a commonly used quantity is the dimensionless energy spectrum density $\Omega_{\rm gw}$. This quantity serves to capture the ratio of GW energy density ${\rm d}\rho_{\rm gw}$ within a specific frequency range [$f$,$f+{\rm d}f$] to the critical energy density $\rho_{\rm c}$, exhibiting a direct relationship with $S_{\rm h}$ Allen (1996); Thrane and Romano (2013):
@@ III.3 Detection sensitivity
The data $s_{I}$ is the linear addition of the SGWB signal $h_{I}$ and the noise $n_{I}$. For channels $I$ and $J$ from two independent detectors, we can define their cross-correlation as Allen and Romano (1999):
$S^{\prime}_{IJ}=\int_{t_{0}-T/2}^{t_{0}+T/2}{\rm d}t\int_{t_{0}-T/2}^{t_{0}+T/2}{\rm d}t^{\prime}s_{I}(t)s_{J}(t^{\prime})Q_{IJ}(t,t^{\prime}),$ (12)
where $Q_{IJ}$ is a filter function that can be determined in the aim to maximize the SNR of the SGWB. Notice though, for reasons that we will explain in detail later, when the noises from channels $I$ and $J$ have non-zero correlation, then the above definition of channel correlation needs to be extended as Smith and Caldwell (2019):
$S_{IJ}=\int_{t_{0}-T/2}^{t_{0}+T/2}{\rm d}t\int_{t_{0}-T/2}^{t_{0}+T/2}{\rm d}t^{\prime}\left[s_{I}(t)s_{J}(t^{\prime})-\langle n_{I}(t)n_{J}(t^{\prime})\rangle\right]Q_{IJ}(t,t^{\prime}),$ (13)
where
@@
For uncorrelated channels, $\langle n_{I}(t)n_{J}(t^{\prime})\rangle=0$, so Eq. (13) reduces to Eq. (12). In real data analysis, the noise PSD needs to be determined through the observation data, which inevitably contains random fluctuation. Fortunately, by utilizing the smooth and stationary property of the noise PSD, we can greatly reduce its random fluctuation by adopting the averaged periodogram method  Welch (1967); Nitz et al. (2018); Dal Canton et al. (2021).
Integrating Eq. (12) into Eq. (13) is a straightforward process that unifies the measurements obtained by the cross-correlation and null-channel methods. In the subsequent analysis, we will examine the SGWB detection using the unified measurement $S_{IJ}$ for both methods. To further streamline subsequent analysis, we designate the detectors as $\mathcal{A}$ and $\mathcal{B}$. Assuming that the employed channels in each detector share the same auto-correlation ORF $\mathcal{R}_{\mathcal{A}}$ and noise PSD, the total SNR in the weak-signal limit Seto (2020); Liang et al. (2022)
@@
the SNR threshold $\rho_{\rm thr}$, beyond which the GW detectors can find the SGWB, can be translated into the threshold of SGWB strength $\Omega_{0}$ based on Eq. (15):
@@ IV Beyond weak-signal limit
Assuming that all other conditions specified in Eq. (13) hold true, except for the weak-signal limit. Then for all channel pairs $\{IJ\}$ (or equivalently, $\{MN\}$), the expectation and variance of the detection measurement
@@
where the filter function $\widetilde{Q}_{IJ}$ is arbitrary. Furthermore, for the last line of Eq. (26) to be valid, it is necessary that the ORF for cross-correlation is significantly lower than that for auto-correlation, leading to
@@
From Eq. (28), it is evident that the extension from Eq. (12) to Eq. (13) guarantees an SNR of 0 in the absence of an SGWB signal.
@@
$\rho_{\rm opt}^{2}=\frac{2}{1+\delta_{\mathcal{A}\mathcal{B}}}\frac{\big{(}\widetilde{Q}_{IJ}(f,t_{0}),\frac{f^{3}\Gamma^{*}_{IJ}(f,t_{0})\Omega_{\rm gw}(f)}{D_{\mathcal{A}\mathcal{B}}(f)}\big{)}^{2}}{\big{(}\widetilde{Q}_{IJ}(f,t_{0}),\widetilde{Q}_{IJ}(f,t_{0})\big{)}}.$ (30)
@@
$\widetilde{Q}_{IJ}(f,t_{0})=\lambda\frac{f^{3}\Gamma^{*}_{IJ}(f,t_{0})\Omega_{\rm gw}(f)}{D_{\mathcal{A}\mathcal{B}}(f)},$ (31)
@@
In contrast to Eq. (15), Eq. (32) suggests that the SGWB detection can suffer negative impacts from the SGWB itself, especially in cases of loud SGWB. To visualize this impact, plotting sensitivity curves is beneficial. Although a direct derivation of $\Omega_{0}$ and $\Omega_{\rm PIS}$ directly from Eq. (19) and Eq. (23) based on Eq. (32) to generate the sensitivity curves proves challenging, these values can still be determined by inverting the SNR formula [^1] In this work, we utilize the Newton-Raphson method.. Fig. 3 and Fig. 4 illustrate the PLIS and PIS curves correspondingly, with solid lines in red, green, and blue denoting TianQin, the TianQin I+II network, and the TianQin + LISA network  [^2] Note that for the TianQin detector, the auto-correlations within the same channel are considered. In the case of detector networks, the cross-correlations between channels from different detectors are taken into consideration., surpassing the weak-signal limit. Dashed lines including the weak-signal limit are provided for comparison. As an illustrative example, the spectral function for the PIS curves is determined by
@@ V Case study
Furthermore, it has been highlighted that the SGWB originating from extragalactic double white dwarfs may exhibit significant intensity levels, leading to high SNR Liang et al. (2022). Without loss of generality, the corresponding energy spectrum density of SGWB can be described as the cumulative contribution from each binary system Phinney (2001); Abbott et al. (2016); Romano and Cornish (2017); Abbott et al. (2018):
@@
where the population parameter $\bm{\theta}$ primarily encompasses the component masses of the system. Given the rarity of astrophysical compact stars at high redshifts due to the lack of star formation, the Hubble parameter $H(z)$ can be accurately approximated as $H_{0}\sqrt{\Omega_{\rm m}(1+z)^{3}+\Omega_{\Lambda}}$  Aghanim et al. (2020). In this paper, we focus on the DA and DB white dwarfs, for which the mass distribution can be modeled by a Gaussian distribution with mean $\mu=0.663\,\,M_{\odot}$ and standard deviation $\sigma=0.177\,\,M_{\odot}$ Rosado (2011). The merger rate is directly linked to the redshift and can be expressed as follows Abbott et al. (2023):
@@
with the local merger rate $R_{0}$ and index $\kappa$. As suggested in Ref. Rosado (2011), the $R_{0}$ is assumed to be within the range of 20 to $500\,\,{\rm Mpc^{-3}\,Myr^{-1}}$. The energy spectrum emitted in the source frame, denoted by ${\rm d}E_{\rm gw}/{\rm d}f_{\rm s}$, is characterized by the frequency in the source frame $f_{\rm s}=f(1+z)$ Zhu et al. (2013). Given that the radius of a white dwarf significantly exceeds its Schwarzschild radius, it is reasonable to assume that below the cut-off frequency where both stars come into contact, the binary systems consistently remain in the inspiral phase.
@@
Regarding the first-order PTs [^3] Unless otherwise specified, we adopt the natural unit for discussions involving first-order PTs., we adopt the model mentioned in our previous study Liang et al. (2022), where the energy spectrum density is formulated as
@@
$\tilde{\Omega}_{i}(\bm{\theta})=\,\,$ $1.67\times 10^{-5}\left(\frac{100}{g_{*}(T_{*})}\right)^{1/3}\left(\frac{H_{*}}{\beta}\right)^{p}$ (39)
×$\times\left(K_{i}(\alpha)\frac{\alpha}{1+\alpha}\right)^{q}\tilde{\Delta}_{i*}(v_{w}),$
@@
Figure. 5 illustrates the SGWBs from various sources: the purple line represents the Galactic foreground with a one-year operation time, the gray region corresponds to the extragalactic background associated with the local merger rate ranging from 20 to $500\,\,{\rm Mpc^{-3}\,Myr^{-1}}$, and the orange line signifies the cosmological background determined by the parameters $\{\alpha,\alpha_{\infty},v_{w},\beta/H_{*},T_{*}\}=\{2,1,0.95,50,10^{3}\,{\rm{GeV}}\}$. To facilitate comparison between the SGWB intensity and noise levels, the noise energy spectrum density of TianQin, the TianQin I+II network, and the TianQin + LISA network are shown in red, green, and blue lines, respectively. While the extragalactic background can near the detector noise level, both the Galactic foreground and cosmological background can exceed the noise. Taking the Galactic foreground as an example, Table. 3 presents the detection SNR within different detector configurations. It also provides the SNR under weak-signal conditions for the sake of comparison. In the weak-signal limit, the TianQin + LISA network achieves the highest SNR for detecting the Galactic foreground. As moving beyond the weak-signal limit, the SNR drops to approximately 20% for the TianQin + LISA network, whereas it remains around 60% for TianQin and the TianQin I+II network. Thus, TianQin demonstrates superior performance in terms of SNR.
@@
In Fig. 6, we present SNR contour plots for the extragalactic background, showcasing scenarios with and without the weak-signal limit in the left and right panels for comparison. Not surprisingly, an SNR threshold of 100 leads to significant variations in the detectable parameter regions between the two scenarios, especially for the TianQin + LISA network. Failure to achieve an SNR above 100 within one year of operation would shift the lower limit for $\kappa$ from around 0 to 2.5, indicating a transition from a scenario with the weak-signal limit to one without it. This shift suggests that the adverse effects of loud SGWB self-generation on detection reduce the region of detectable parameters to a lesser extent than initially anticipated. On the other hand, assuming a constant merger rate with redshift ($\kappa=0$), achieving an SNR exceeding 100 after 1 year of operation, under the weak-signal limit, would estimate the merger rate for EDWDs to be around $500\,\,{\rm Mpc^{-3}\,Myr^{-1}}$. However, this value falls considerably short of meeting the actual detection requirements.
A similar diminishing effect of parametric limitations can be observed in Fig. 7, where we follow Refs. Caprini et al. (2016, 2020) and set the parameters $\{\alpha,\alpha_{\infty},v_{w}\}=\{2,1,0.95\}$ for the first-order PTs. By setting $\alpha$ greater than $\alpha_{\infty}$, the SGWB is anticipated to involve the collision, acoustic, and turbulent phases; the bubble velocity $v_{w}$ exceeding the sound speed $c_{\rm s}$ corresponds to detonation bubble collisions Kamionkowski et al. (1994). Note that, changing other PT parameters and the parameter degeneracy may also influence the results depicted in the Fig. 7. If the SNR surpasses 100 after 1 year of operation, the temperature $T_{*}$ is expected to be significantly greater than $10\,\,{\rm{GeV}}$, while the PT duration $\beta/H_{*}$ should be less than $10^{2}$. The detectable region for these two parameters can expand notably by utilizing the weak-signal limit, leading to potential misjudgments in detection. It is important to highlight that, due to the wide range of variability in fixed parameters, effectively constraining the parameters of PTs remains a challenging task.
Figure 6: SNR contour plots for the SGWB resulting from extragalactic DWD. The left-hand and right hand panels refer to the scenarios with and without weak-signal limit, considering a one-year operation time.